# Friedrich-Hirzebruch-Allee 5

### v1

@@ Abstract
Autonomous systems, such as self-driving cars, rely on reliable semantic environment perception for decision making. Despite great advances in video semantic segmentation, existing approaches ignore important inductive biases and lack structured and interpretable internal representations. In this work, we propose MCDS-VSS, a structured filter model that learns in a self-supervised manner to estimate scene geometry and ego-motion of the camera, while also estimating the motion of external objects. Our model leverages these representations to improve the temporal consistency of semantic segmentation without sacrificing segmentation accuracy. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate motion of dynamic objects, and finally the predicted scene features are fused with the current features to obtain a temporally consistent scene segmentation. Our model parses automotive scenes into multiple decoupled interpretable representations such as scene geometry, ego-motion, and object motion. Quantitative evaluation shows that MCDS-VSS achieves superior temporal consistency on video sequences while retaining competitive segmentation performance. Animations and code to reproduce our results can be found in our project website.
@@ 1 Introduction
Through self-supervised learning (SSL), MCDS-VSS learns to estimate scene geometry and ego-motion. It also estimates motion of additional objects (e.g. pedestrians or vehicles), and hard-wires our knowledge from the MCDS domain to project the previous scene features into the current time-step using these representations. The structured design of our filter allows us to factorize the perceived complex changes in the scene into simpler factors of variation; thus easing the modeling of temporal information.
@@ 2 Related Work
VSS methods are often divided into two distinct categories. The first class aims to reduce the computational cost and improve the efficiency of segmentation models, instead of naively encoding and interpreting every single input frame. Several methods improve the efficiency by propagating and reusing features extracted from selected key frames [Jain et al.(2019)Jain, Wang, and Gonzalez, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei]; whereas other approaches achieve efficiency by employing lightweight neural network blocks [Pfeuffer and Dietmayer(2019), Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell] or by distilling the information from large teacher models to smaller models [Liu et al.(2020)Liu, Shen, Yu, and Wang]. The second category, to which our proposed method belongs, aims to improve the semantic segmentation performance and temporal consistency by exploiting the temporal continuity of video streams. Some methods exploit temporal dependencies between video frames and improve the consistency of the predicted segmentation maps by combining image segmentation models with recurrent neural networks (RNNs) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer, Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell, Villar-Corrales et al.(2022)Villar-Corrales, Karapetyan, Boltres, and Behnke] or with attention-based modules [Wang et al.(2021a)Wang, Wang, and Liu, Lao et al.(2023)Lao, Hong, Guo, Zhang, Wang, Chen, and Chu, Sun et al.(2024)Sun, Liu, Ding, Wu, and Van Gool, Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang, Sun et al.(2022)Sun, Liu, Tang, Chhatkuli, Zhang, and Van Gool]. Another family of works use an optical flow module to compute the feature correspondence between consecutive frames, and then use this flow for predictive feature learning [Ding et al.(2020)Ding, Wang, Zhou, Shi, Lu, and Luo, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei, Liu et al.(2020)Liu, Shen, Yu, and Wang, Xu et al.(2018)Xu, Fu, Yang, and Lee, Baghbaderani et al.(2024)Baghbaderani, Li, Wang, and Qi, Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt, Hur and Roth(2016)].
Our method belongs to the latter category. However, unlike aforementioned approaches, MCDS-VSS incorporates assumptions from the domain of moving cameras and dynamic scenes into the model design, computing interpretable intermediate representations, which lead to more temporally consistent video segmentation results.
@@
Several works have investigated the use of depth and motion for semantic segmentation in videos. Approaches like [Shi et al.(2020)Shi, Lin, Wang, Hung, and Wang, Athar et al.(2023)Athar, Li, Casas, and Urtasun, Cao and Behnke(2024), Zhu et al.(2023)Zhu, Han, Cai, Borse, Ghaffari, and Porikli] segment dynamic scenes by jointly processing video frames with depth information captured by LiDAR scanners; whereas methods like [Ma et al.(2017)Ma, Stückler, Kerl, and Cremers, Kundu et al.(2020)Kundu, Yin, Fathi, Ross, Brewington, Funkhouser, and Pantofaru, Chen et al.(2021)Chen, Lin, Wang, Yang, and Cheng, Rosu et al.(2020)Rosu, Quenzel, and Behnke] use depth and pose information to enforce consistency between segmentation maps from multiple viewpoints. Recently, depth-aware panoptic segmentation (PS) models [Qiao et al.(2021)Qiao, Zhu, Adam, Yuille, and Chen, Petrovai and Nedevschi(2023), Yuan et al.(2022)Yuan, Li, Yang, Cheng, Zhang, Tong, Zhang, and Tao] aim to jointly solve the tasks of PS and depth estimation by extending a segmentation model with a depth decoder and conditioning its prediction using instance-masks.
[…]
@@ 3 MCDS-VSS Structured Filtering Method
We propose MCDS-VSS, illustrated in Fig. 1, a structured filter that improves the temporal consistency of a semantic segmentation model on moving camera dynamic scene scenarios. MCDS-VSS learns in a self-supervised way to estimate geometry and motion representations, i.e., scene depth and camera ego-motion. It also estimates the motion of other agents in the scene, and uses these human-interpretable representations to propagate abstract scene features over time, thus improving its segmentation performance and temporal consistency.
@@ 3.1 Learning of Geometry & Motion
MCDS-VSS learns in a self-supervised manner to estimate scene geometry and camera motion, which are then used to improve the temporal consistency of a segmentation model. Fig. 2 illustrates our two-step self-supervised approach for learning the scene geometry with camera motion and for distillation of object dynamics.
@@
To predict the scene geometry, MCDS-VSS incorporates a depth decoder $\mathcal{D}_{\textrm{d}}$ that outputs the depth $\mathbf{d}_{t}$ and inverse depth $\mathbf{d}^{\triangleleft}_{t}$111 With slight abuse of notation, we denote the $\mathbf{d}^{\triangleleft}_{t}$ as disparities. of the scene given the input feature maps $\mathbf{h}_{t}$; whereas to compute the camera motion between two images we employ a motion encoder $\mathcal{E}_{\textrm{m}}$ that computes motion features between two sets of feature maps, and an ego-motion decoder $\mathcal{D}_{\textrm{c}}$, which predicts the camera transformation between two time steps $\textrm{C}_{{t-1}}^{\,{t}}$, parameterized as a 6-dim. vector containing the translation and Euler angles of the camera transformation matrix. We then render the ego-warped image $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ using the estimated depth and ego-motion:
@@
where $K\in{\mathbb{R}}^{3,3}$ are the camera intrinsic parameters, and $\mathcal{F}_{\textrm{fwd}}$ is the forward rendering function proposed in [Lee et al.(2021)Lee, Im, Lin, and Kweon]. If the depth and ego-motion estimates are accurate, the resulting warped images should match the target images except for occluded regions and moving objects. Therefore, as illustrated in Fig. 2(a), we train the modules for self-supervised depth and ego-motion estimation by optimizing the following loss function:
@@
Figure 2: Learning geometry and motion. a) We learn the depth $\mathbf{d}_{t-1}$ and ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ given two video frames by enforcing a photometric loss between warped and target frame $\mathcal{L}_{\textrm{Photo}}$ and a depth regularization $\mathcal{L}_{\textrm{Reg}}$. b) Given an ego-warped image, we train a flow decoder to model the dynamics of moving objects by distilling a pretrained RAFT model.
@@
As illustrated in Fig. 2(b), we estimate the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ between the ego-warped and target images, which parameterizes the dynamics of moving objects, by training an residual flow decoder $\mathcal{R}_{\textrm{f}}$ while keeping all other modules frozen. $\mathcal{R}_{\textrm{f}}$ is trained to match the optical flow predictions of the large state-of-the-art optical flow model RAFT [Teed and Deng(2020)]:
@@ 3.2 Structured Filter
form the core of the MCDS-VSS structured filter, which is depicted in Fig. 1. It propagates information over time using two different filter states, namely a scene state $\mathbf{s}$ that encodes the scene contents and geometry and a camera state $\mathbf{c}$, which encodes the ego-motion of the vehicle.
@@
Ego-Motion Filter: The ego-motion filter extends the motion encoder and camera predictor modules in order to aggregate motion information over time and enforce the prediction of temporally consistent ego-motion. The temporal integration is achieved via a motion update module, which is implemented as a ConvGRU [Ballas et al.(2016)Ballas, Yao, Pal, and Courville] recurrent layer that jointly processes the motion features and previous camera state $\mathbf{c}_{t-1}$ and outputs an updated camera state $\mathbf{c}_{t}$ from which the ego-motion can be then predicted:
@@
Depth Estimation and Ego-Motion Compensation: Given the past scene state $\mathbf{s}_{t-1}$, the depth decoder $\mathcal{D}_{\textrm{d}}$ computes the depth map $\mathbf{d}_{t-1}$. This scene geometry and the ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ are used as in Eq. (1) to project $\mathbf{s}_{t-1}$ to time $t$. The resulting ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ encodes scene contents and geometry after compensation for ego-motion.
Object-Motion Compensation: This module models dynamic objects in the scene, such as pedestrians or vehicles, and updates the scene state to compensate for the motion of such objects. We jointly process the ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ and the current image features $\mathbf{h}_{t}$ with the residual flow decoder $\mathcal{R}_{\textrm{f}}$ in order to compute the residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$, which represents the pixel displacement of moving objects between consecutive time steps. The ego-warped features $\mathbf{s}^{\textrm{ego}}_{t}$ are then projected into the current time-step by applying the displacement encoded in the residual flow map, followed by bilinear interpolation to obtain valid coordinate values. The resulting features $\mathbf{s}^{\textrm{full}}_{t}$ not only incorporate the motion of dynamic objects in the scene, but can also correct alignment errors between $\mathbf{s}^{\textrm{ego}}_{t}$ and $\mathbf{h}_{t}$ that might occur due to inaccurate depth or ego-motion estimates.
@@ 4.1 Model Training
To ease the training process, we propose a multi-stage approach in which we first train the encoder and decoder modules using image pairs or triplets, and then integrate and train the filter modules using sequences of six frames in order to improve segmentation performance and temporal consistency while retaining interpretable representations.
Table 1: Training stages and hyper-parameters.
@@
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$   (2) $2\cdot 10^{-4}$ 3
@@ 3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$   (9) $8\cdot 10^{-5}$ 6
4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$   (10) $8\cdot 10^{-5}$ 6
@@
, MCDS-VSS encoder and decoder modules are jointly trained for self-supervised learning of geometry and ego-motion, as well as for semantic segmentation by minimizing a combination of cross entropy $\mathcal{L}_{\textrm{Seg}}$ and SSL geometry $\mathcal{L}_{\textrm{Depth}}$ losses. Following [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], we use image triplets $(\mathbf{x}_{t-\tau},\mathbf{x}_{t},\mathbf{x}_{t+\tau})$, with $\mathbf{x}_{t}$ being the target. Subsequently, we train the residual flow decoder as described in
@@
while keeping the remaining modules frozen. In the third stage, with the goal of improving the temporal continuity of the predicted ego-motion, we train the ego-motion filter and compensation by minimizing the loss function:
@@
which enforces the model to compute accurate camera motion estimates in order to align the ego-warped state with the current observations. Finally, we jointly train the fusion module and fine-tune the segmentation decoder by minimizing the following loss function:
$\mathcal{L}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Seg}}({\mathbf{\hat{y}}_{t}},{\mathbf{y}_{t}})+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}({\mathbf{\tilde{y}}_{t}},{\mathbf{\hat{y}}_{t}}),$ (10)
where $\mathcal{L}_{\textrm{Seg}}$ is the cross entropy loss function and $\mathcal{L}_{\textrm{TC}}$ is a temporal consistency regularizer that enforces the segmentation $\mathbf{\tilde{y}}_{t}$ computed by decoding $\mathbf{s}^{\textrm{full}}_{t}$ to be close to the actual predicted segmentation maps $\mathbf{\hat{y}}_{t}$.

### v2

# Center for Robotics and

@@ Abstract
Autonomous systems, such as self-driving cars, rely on reliable semantic environment perception for decision making. Despite great advances in video semantic segmentation, existing approaches ignore important inductive biases and lack structured and interpretable internal representations. In this work, we propose MCDS-VSS, a structured filter model that learns in a self-supervised manner to estimate scene geometry and ego-motion of the camera, while also estimating the motion of external objects. Our model leverages these representations to improve the temporal consistency of semantic segmentation without sacrificing segmentation accuracy. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate motion of dynamic objects, and finally the predicted scene features are fused with the current features to obtain a temporally consistent scene segmentation. Our model parses automotive scenes into multiple decoupled interpretable representations such as scene geometry, ego-motion, and object motion. Quantitative evaluation shows that MCDS-VSS achieves superior temporal consistency on video sequences while retaining competitive segmentation performance. Code and pretrained models are available in the project website.
@@ 1 Introduction
Through self-supervised learning (SSL), MCDS-VSS learns to estimate scene geometry and ego-motion. It also estimates motion of additional moving objects (e.g. pedestrians or vehicles), and hard-wires our knowledge from the MCDS domain to project the previous scene features into the current time-step using these representations. The structured design of our filter allows us to factorize the perceived complex changes in the scene into simpler factors of variation; thus easing the modeling of temporal information.
@@ 2 Related Work
VSS methods are often divided into two distinct categories. The first class aims to reduce the computational cost and improve the efficiency of segmentation models, instead of naively encoding and interpreting every single input frame. Several methods improve the efficiency by propagating and reusing features extracted from selected key frames [Jain et al.(2019)Jain, Wang, and Gonzalez, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei]; whereas other approaches achieve efficiency by employing lightweight neural network blocks [Pfeuffer and Dietmayer(2019), Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell] or by distilling the information from large teacher models into smaller models [Liu et al.(2020)Liu, Shen, Yu, and Wang]. The second category, to which our proposed method belongs, aims to improve the semantic segmentation performance and temporal consistency by exploiting the temporal continuity of video streams. Some methods exploit temporal dependencies between video frames and improve the consistency of the predicted segmentation maps by combining image segmentation models with recurrent neural networks (RNNs) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer, Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell, Villar-Corrales et al.(2022)Villar-Corrales, Karapetyan, Boltres, and Behnke] or with attention-based modules [Wang et al.(2021a)Wang, Wang, and Liu, Lao et al.(2023)Lao, Hong, Guo, Zhang, Wang, Chen, and Chu, Sun et al.(2024)Sun, Liu, Ding, Wu, and Van Gool, Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang, Sun et al.(2022)Sun, Liu, Tang, Chhatkuli, Zhang, and Van Gool]. Another family of works use an optical flow module to compute the feature correspondence between consecutive frames, and then use this flow for predictive feature learning [Ding et al.(2020)Ding, Wang, Zhou, Shi, Lu, and Luo, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei, Liu et al.(2020)Liu, Shen, Yu, and Wang, Xu et al.(2018)Xu, Fu, Yang, and Lee, Baghbaderani et al.(2024)Baghbaderani, Li, Wang, and Qi, Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt, Hur and Roth(2016)].
Our method belongs to the latter category of VSS models. However, unlike aforementioned approaches, MCDS-VSS incorporates assumptions from the domain of moving cameras and dynamic scenes into the model design, and computes interpretable intermediate geometry and motion-aware representations, which lead to accurate and temporally consistent video segmentation results.
@@
Several works have investigated the use of depth and motion for semantic segmentation in videos. Approaches like [Shi et al.(2020)Shi, Lin, Wang, Hung, and Wang, Athar et al.(2023)Athar, Li, Casas, and Urtasun, Cao and Behnke(2024), Zhu et al.(2023)Zhu, Han, Cai, Borse, Ghaffari, and Porikli] segment dynamic scenes by jointly processing video frames with depth information captured by LiDAR scanners; whereas methods like [Ma et al.(2017)Ma, Stückler, Kerl, and Cremers, Kundu et al.(2020)Kundu, Yin, Fathi, Ross, Brewington, Funkhouser, and Pantofaru, Chen et al.(2021)Chen, Lin, Wang, Yang, and Cheng, Rosu et al.(2020)Rosu, Quenzel, and Behnke] use depth and camera pose information in combination with a semantic segmentation model in order to improve the segmentation performance by enforcing consistency between predictions from multiple viewpoints. Recently, depth-aware panoptic segmentation models [Qiao et al.(2021)Qiao, Zhu, Adam, Yuille, and Chen, Petrovai and Nedevschi(2023), Yuan et al.(2022)Yuan, Li, Yang, Cheng, Zhang, Tong, Zhang, and Tao] aim to jointly solve the tasks of panoptic segmentation and depth estimation by extending a segmentation model with a depth decoder and conditioning its prediction using instance-masks.
@@ 3 MCDS-VSS Structured Filtering Method
We propose MCDS-VSS, illustrated in Figure 1, a structured filter that improves the temporal consistency of a semantic segmentation model on moving camera dynamic scene scenarios. MCDS-VSS learns in a self-supervised way to estimate geometry and motion representations, i.e., scene depth and camera ego-motion. It also estimates the motion of other agents in the scene, and uses these human-interpretable representations to propagate abstract scene features over time, thus improving its segmentation performance and temporal consistency.
@@ 3.1 Learning of Geometry & Motion
MCDS-VSS learns in a self-supervised manner to estimate scene geometry and camera motion, which are then used to improve the temporal consistency of a segmentation model. Figure 2 illustrates our two-step self-supervised approach for learning the scene geometry with camera motion and for distillation of object dynamics.
@@
To predict the scene geometry, MCDS-VSS incorporates a depth decoder $\mathcal{D}_{\textrm{d}}$ that outputs the depth $\mathbf{d}_{t}$ and inverse depth $\mathbf{d}^{\triangleleft}_{t}$111 With slight abuse of notation, we denote the $\mathbf{d}^{\triangleleft}_{t}$ as disparities. of the scene given the input feature maps $\mathbf{h}_{t}$; whereas to compute the camera motion between two images we employ a motion encoder $\mathcal{E}_{\textrm{m}}$ that computes motion features between two sets of feature maps, and an ego-motion decoder $\mathcal{D}_{\textrm{c}}$, which predicts the camera transformation between two time steps $\textrm{C}_{{t-1}}^{\,{t}}$, parameterized as a 6-dimensional vector containing the translation parameters and Euler angles of the camera transformation matrix. We then render the ego-warped image $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ using the estimated scene depth and ego-motion:
@@
where $K\in{\mathbb{R}}^{3,3}$ are the camera intrinsic parameters, and $\mathcal{F}_{\textrm{fwd}}$ is the forward rendering function proposed in [Lee et al.(2021)Lee, Im, Lin, and Kweon]. If the depth and ego-motion estimates are accurate, the resulting warped images should match the target images except for occluded regions and moving objects. Therefore, as illustrated in Figure 2(a), we train the modules for self-supervised depth and ego-motion estimation by optimizing the following loss function:
@@
Figure 2: Learning geometry and motion. a) We learn the scene depth $\mathbf{d}_{t-1}$ and ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ in a self-supervised manner given two video frames by enforcing a photometric loss $\mathcal{L}_{\textrm{Photo}}$ between the ego-warped $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ and target frames $\mathbf{x}_{t}$, as well as a depth regularization $\mathcal{L}_{\textrm{Reg}}$. b) Given an ego-warped image, we train a residual flow decoder to predict the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ that parameterizes the dynamics of moving objects in the scene by distilling a pretrained RAFT model.
@@
As illustrated in Figure 2(b), we estimate the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ between the ego-warped and target images, which encodes the dynamics of moving objects, by training a residual flow decoder $\mathcal{R}_{\textrm{f}}$ while keeping all other modules frozen. The residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ is parameterized as a 2D flow field that encodes the per-pixel motion in the horizontal and vertical directions needed to align the ego-warped images $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ to the corresponding target images $\mathbf{x}_{t}$. $\mathcal{R}_{\textrm{f}}$ is trained to match the optical flow predictions of the large state-of-the-art optical flow model RAFT [Teed and Deng(2020)]:
@@ 3.2 Structured Filter
form the core of the MCDS-VSS structured filter, which is depicted in Figure 1. It propagates information over time using two different filter states, namely a scene state $\mathbf{s}$ that encodes the scene contents and geometry, and a camera state $\mathbf{c}$ that encodes the ego-motion of the vehicle.
@@
Ego-Motion Filter: The ego-motion filter extends the motion encoder $\mathcal{E}_{\textrm{m}}$ and ego-motion decoder $\mathcal{D}_{\textrm{c}}$ modules in order to aggregate motion information over time and enforce the prediction of temporally consistent ego-motion. The temporal integration is achieved via a motion update module, which is implemented as a ConvGRU [Ballas et al.(2016)Ballas, Yao, Pal, and Courville] recurrent layer that jointly processes the motion features and previous camera state $\mathbf{c}_{t-1}$, and outputs an updated camera state $\mathbf{c}_{t}$ from which the ego-motion can be then predicted:
@@
Depth Estimation and Ego-Motion Compensation:. Given the past scene state $\mathbf{s}_{t-1}$, the depth decoder $\mathcal{D}_{\textrm{d}}$ computes the depth map $\mathbf{d}_{t-1}$. This scene geometry and the estimated ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ are used as in Equation (1) to project $\mathbf{s}_{t-1}$ to time $t$. The resulting ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ encodes scene contents and geometry after compensation for ego-motion.
Residual Flow Estimation and Object Motion Compensation: These modules model dynamic objects in the scene, such as pedestrians or vehicles, and update the scene state to compensate for the motion of such objects. We jointly process the ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ and the current image features $\mathbf{h}_{t}$ with the residual flow decoder $\mathcal{R}_{\textrm{f}}$ in order to compute the residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$, which represents the pixel displacement of moving objects between consecutive time steps. The ego-warped features $\mathbf{s}^{\textrm{ego}}_{t}$ are then projected into the current time-step by applying the displacement encoded in the residual flow map, followed by bilinear interpolation to obtain valid coordinate values. The resulting features $\mathbf{s}^{\textrm{full}}_{t}$ not only incorporate the motion of dynamic objects in the scene, but can also correct alignment errors between $\mathbf{s}^{\textrm{ego}}_{t}$ and $\mathbf{h}_{t}$ that might occur due to inaccurate depth or ego-motion estimates.
@@ 4.1 Model Training
To ease the training process, we propose a multi-stage training procedure in which we first train the encoder and decoder modules using image pairs or triplets, and then integrate and train the filter modules using sequences of six frames in order to gather scene context information and improve the segmentation performance and temporal consistency while retaining interpretable representations.
Table 1: MCDS-VSS training stages and hyper-parameters.
@@
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$   (2) $2\cdot 10^{-4}$ 3
@@ 3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$   (9) $8\cdot 10^{-5}$ 6
4 Temporal Integration $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$   (10) $8\cdot 10^{-5}$ 6
@@
, MCDS-VSS encoder and decoder modules are jointly trained for self-supervised learning of geometry and ego-motion, as well as for semantic segmentation by minimizing a combination of cross entropy $\mathcal{L}_{\textrm{Segm}}$ and SSL geometry $\mathcal{L}_{\textrm{Depth}}$ losses. Following [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], we use image triplets $(\mathbf{x}_{t-\tau},\mathbf{x}_{t},\mathbf{x}_{t+\tau})$, with $\mathbf{x}_{t}$ being the target image and $\tau$ being the temporal distance between source and target frames used during this first training stage. Subsequently, we train the residual flow decoder using image pairs as described in
@@
while keeping the remaining modules frozen. In the third stage, with the goal of improving the temporal continuity of the predicted ego-motion, we train the ego-motion filter and compensation using short video sequences of length $T$ by minimizing the loss function:
@@
which enforces the model to compute accurate camera motion estimates in order to align the ego-warped state with the current observations. Finally, in the last training stage we jointly train the feature fusion module and fine-tune the segmentation decoder by minimizing the following loss function:
$\mathcal{L}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Segm}}({\mathbf{\hat{y}}_{t}},{\mathbf{y}_{t}})+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}({\mathbf{\tilde{y}}_{t}},{\mathbf{\hat{y}}_{t}}),$ (10)
where $\mathcal{L}_{\textrm{Segm}}$ is the cross entropy loss function and $\mathcal{L}_{\textrm{TC}}$ is a temporal consistency regularizer that enforces the segmentation $\mathbf{\tilde{y}}_{t}$ computed by decoding $\mathbf{s}^{\textrm{full}}_{t}$ to be close to the actual predicted segmentation maps $\mathbf{\hat{y}}_{t}$.